OpenStudy (gigigirl0623):

A graph is shown below: Solid line joining 0, negative 2 and 1, 3 and the region below the line is shaded. Which of the following inequalities is best represented by this graph? 5x + y ≤ 2 5x + y ≥ 2 5x − y ≤ 2 5x − y ≥ 2

jimthompson5910 (jim_thompson5910):

First let's find the slope of the line through (0,-2) and (1,3) Do you know how to do this?

OpenStudy (gigigirl0623):

No

jimthompson5910 (jim_thompson5910):

do you agree that for something like (0,-2) the value of x is 0 and the value of y is -2 ?

OpenStudy (sshayer):

Let us find an equation through (0,-2) and (1,3) \[y-y1=\frac{ y2-y1 }{ x2-x1 }\left( x-x1 \right)\] \[y+2=\frac{ 3-(-2) }{ 1-0 }\left( x-0 \right)\] y+2=5x 5x-y=2 now let us draw this line